At what speed do a bicycle and its rider, with a combined mass of 100 kg, have the same momentum as a 1300kg car traveling at 1.2m/s ?
Answer #1

m1 * v1 = m2 * v2

100 * v = 1300 * 1.2

v = 15.6 m/s
Answer #2


momentum of bicycle and rider is p1 = m1*u1= 100*u1...

momentum of car is p2 = m2*u2 =1300*1.2 = 1560 kgm/s....

given that 1560 = 100*u1.....

u1 = 1560/100 = 15.6 m/sec